You are here: Foswiki>Commissioning Web>Com20080903 (13 Oct 2008, NormCushing)Edit Attach

Commissioning Notes from 2nd-3rd September 2008 MST

Observers: D.Miller, T.Leibold for a couple hours in Tucson, and for the second half J. Storm in Potsdam
OSA: A. Ceranski

Summary

Thunderstorms over Mt Graham and in the vicinity (~100% humidity) are keeping us closed at sunset

AzCam not working in AGw so can not perform any BG Commissioning

Collected long blocks of vibration data with the IRTC and EF data with the IRTC to calibrate EF with primary modes

Night time Activities

Times in MST below

VNC to AZ-Cam computer

VNC connectivity to the AzCam Server

Opened @ 2:05

2:05 Opened to 4 mags extinction. AGw cameras are still not working, but will take IRTC data.

2:50 After troubles with irs (now must be running on lbmtu01and was not), could not see a star on the IRTC. spiral pattern: find star at 90, 30 Aaron reports sky is looking pretty good: haze is gone.

2:58 ptlist: IE -56.27630, CA -37.00000

3:05 0,0 offset, IE -100.00000, CA -125.00000, bunch of coma and focus

3:15 z7 = 3500, z8 = 2500 to remove coma by eye

3:20 Z4 = 5500

3:30 trying to get my IDL getimage routines to work

3:40 Setting up IRTC to save data in /Repository/IRTC_Data/080903

3:50 set IE=-95, CA = -135 to center star on IRTC

3:53 0.003 sec exposure, 5000 frames. This got IRTC in a continuous loop. Take ~ 3200 of 5000 and then start again at 0. Had to kill IRTC control with task manager.

4:10 vibrate_00009_cube.fit 0.001 sec exposure, 3000 frames 55 elevation az 229, wind NW @ 2 meter/s

4:20 vibrate_00010_cube.fit 0.01 sec exposure, 3000 frames 55 elevation az 229, wind NW @ 2 meter/s

4:50 New star M5_0094 IE = -100, CA=-80

4:55 EF data

ef_000*.fit in/out 0.06/-0.6 all data in /Repository/080903

Intra Extra Mode Amplitude
3 4   none
7 6 z4 +1000
8 9 z5 +1000
11 10   none
12 13 z7 +500
15 14 z9 +500
16 17 none
19 18 z11 +500
20 21   none
23 22 z6 +1000
24     background
25 26 z8 +500
28 27 Z10 +500
29 30   none
bending of z22 failed with Bad
32 31 none
collimate button on primary GUI
33 34   none
35     background

-- DougMiller - 03 Sep 2008
Topic revision: r7 - 13 Oct 2008, NormCushing
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback